OpenStudy (anonymous):

Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 3x - 18 = 2x - 20 Equation #2 10 - 2x = 4x + 34 - The solution to equation #1 is smaller than the solution to equation #2. - The solution to equation #1 is larger than the solution to equation #2. - The solution to equation #1 is the opposite of the solution to equation #2. - None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

- The solution to equation #1 is larger than the solution to equation #2.

OpenStudy (anonymous):

solve them both to see \[3x-18=2x-20\] \[3x-2x=-20+18\] \[x=-2\] is the first one second one is \[10-2x=4x+34\] \[10-34=4x+2x\] \[-24=6x\] \[x=-4\] now compare